if the probability of getting a lift to work on any morning is 0.6 .X is the number of days in a working week of 5 days
(a) what is the probability distribution of X? (please justify)
(b) what is the probability that in a working week will get a lift only twice?

Answer by stanbon(75887)

if the probability of getting a lift to work on any morning is 0.6 .X is the number of days in a working week of 5 days
(a) what is the probability distribution of X? (please justify)
mean = np = 5*0.6 = 3
std = sqrt(npq) = sqrt(3*0.4) = sqrt(1.2)
----------------------------------
(b) what is the probability that in a working week will get a lift only twice?
P(x = 2) = 5C2*0.6^2*0.4^3 = binompdf(5,0.6,2) = 0.2304
------------
Cheers,
Stan H.
--------